# The entropy of dynamical de Sitter horizons

### v3

@@ Abstract
Recently Hollands, Wald and Zhang proposed a new formula for the entropy of a dynamical black hole. We lift this construction to the cosmological event horizon of an asymptotically de Sitter spacetime. By introducing a nontrivial correction term, we generalize Gibbons and Hawking’s first law of event horizons to non-stationary eras. We also develop the non-stationary physical process first law between two arbitrary horizon cross-sections for the cosmological event horizon.
@@ 1 Introduction
Shortly after the discovery of black hole thermodynamics, Gibbons and Hawking established the thermodynamics for the “cosmological event horizons” of de Sitter spacetimes [6]. One of the key results they found is the “first law of event horizons”, which states that the variation away from a Kerr-de Sitter black hole spacetime satisfies
@@ 2 Review of the dynamical black hole entropy
where $k^{a}$ and $l^{a}$ denote the (future directed) outgoing and ingoing null normal to the cross-section respectively, and they are normalized as $k^{a}l_{a}=-1$. $\boldsymbol{\epsilon}_{\mathcal{C}}$ represents the codimension22-2- 2 form of the cross-section. With suitable gauge conditions on perturbations one can prove that[5, 4]
$\boldsymbol{Q}[\xi]\overset{\mathcal{C}}{=}-\frac{1}{16\pi G}\boldsymbol{\epsilon}_{\mathcal{C}}(k_{a}l_{b}-l_{a}k_{b})\nabla^{a}\xi^{b}\overset{\mathcal{C}}{=}\frac{\kappa}{8\pi G}\boldsymbol{\epsilon}_{\mathcal{C}},$ (2.25)
@@ 3 Geometric setup
Consider an electrically neutral asymptotically de Sitter stationary black hole spacetime as shown in Figure 2. We shall adopt the definition of [6] to define the event horizon of the spacetime as the boundary of the past of $\lambda$, i.e., $\overset{\cdot}{I}(\lambda)$, where $\lambda$ is a future inextensible timelike curve representing an observer’s world line. We assume that both the black hole event horizon $\mathcal{H}_{H}$ and the cosmological event horizon $\mathcal{H}_{C}$ possess the structure of bifurcate Killing horizons. And we are mainly interested in the parts of horizons that lie to the future of the bifurcation surfaces $\mathcal{B}_{H}$ and $\mathcal{B}_{C}$. We denote the affinely parameterized null generators of the black hole event horizon and the cosmological event horizon by $k_{H}^{a}$ and $k_{C}^{a}$ respectively, and set the affine parameters to be 00 at the bifurcation surfaces. The Killing field that is normal to the cosmological event horizon is denoted by $\xi^{a}$. Then the Killing vector which coincides with the generators of the black hole event horizon can be expressed in the form[6]
@@ 4 The non-stationary first law for de Sitter horizons
In this section we utilize the covariant phase space formalism to derive both the non-stationary comparison first law for de Sitter black holes, and the physical process first law for cosmological event horizons.
@@ 4.1 The comparison first law
Since for a stationary de Sitter black hole, $J_{H}=\int_{\mathcal{C}_{H}}\boldsymbol{Q}_{g}[\varphi]=\frac{1}{16\pi}\int_{H}\boldsymbol{\epsilon}_{ab}\nabla^{a}\varphi^{b}$ recovers the definition of the angular momentum of the black hole[6]. Therefore the non-stationary comparison first law between an arbitrary black hole horizon cross-section $\mathcal{C}_{H}$ and an arbitrary cosmological horizon cross-section $\mathcal{C}_{C}$ reads
@@ 4.2 The physical process first law
Figure 3: The physical process first law for the de Sitter horizon. An external matter source $\delta T_{ab}$ passes through the cosmological event horizon between two arbitrary cross-sections $\mathcal{C}_{C}(v_{1})$ and $\mathcal{C}_{C}(v_{2})$.
@@
$\Delta\delta M=-\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}(v)}\mathrm{d}A\ \delta T_{ab}\xi^{a}k^{b},$ (4.17)
@@ 5 Conclusion and discussion
We have lifted the construction of the dynamical entropy to cosmological event horizons in asymptotically de Sitter spacetimes. By applying the Noether charge method to non-stationary perturbations of a stationary de Sitter black hole spacetime, we have demonstrated that our formula satisfies both the non-stationary comparison first law and the non-stationary physical process first law. All of these further encourage us to extend the dynamical entropy to more general horizons[22, 23], such as expansion and shear free horizons.

### v4

@@ Abstract
Recently Hollands, Wald and Zhang proposed a new formula for the entropy of a dynamical black hole. We lift this construction to the dynamical cosmological event horizon of an asymptotically de Sitter spacetime. By introducing a nontrivial correction term in the formula for the entropy, we generalize Gibbons and Hawking’s “first law of event horizons” to non-stationary eras. We also develop the non-stationary physical process first law between two arbitrary horizon cross-sections for the cosmological event horizon.
@@ 1 Introduction
Shortly after the discovery of black hole thermodynamics, Gibbons and Hawking established the thermodynamics for the cosmological event horizons of de Sitter spacetimes [6]. One of the key results they found is the “first law of event horizons”, which states that the variation away from a Kerr-de Sitter black hole spacetime satisfies
@@ 2 Review of the dynamical black hole entropy
where $k^{a}$ and $l^{a}$ denote the (future directed) outgoing and ingoing null normal to the cross-section respectively, and they are normalized as $k^{a}l_{a}=-1$. $\boldsymbol{\epsilon}_{\mathcal{C}}$ represents the codimension22-2- 2 form of the cross-section. So that
$\boldsymbol{Q}[\xi]\overset{\mathcal{C}}{=}-\frac{1}{16\pi G}\boldsymbol{\epsilon}_{\mathcal{C}}(k_{a}l_{b}-l_{a}k_{b})\nabla^{a}\xi^{b}\overset{\mathcal{C}}{=}\frac{\kappa}{8\pi G}\boldsymbol{\epsilon}_{\mathcal{C}}.$ (2.25)
And with suitable gauge conditions on perturbations one can prove that[5, 4]
@@ 3 Geometric setup
Consider an electrically neutral asymptotically de Sitter stationary black hole spacetime as shown in Figure 2. We shall adopt the definition of [6] to define the event horizon of the spacetime as the boundary of the past of $\lambda$, i.e., $\dot{I}(\lambda)$, where $\lambda$ is a future inextensible timelike curve representing an observer’s world line. We assume that both the black hole event horizon $\mathcal{H}_{H}$ and the cosmological event horizon $\mathcal{H}_{C}$ possess the structure of bifurcate Killing horizons. And we are mainly interested in the parts of horizons that lie to the future of the bifurcation surfaces $\mathcal{B}_{H}$ and $\mathcal{B}_{C}$. We denote the affinely parameterized null generators of the black hole event horizon and the cosmological event horizon by $k_{H}^{a}$ and $k_{C}^{a}$ respectively, and set the affine parameters to be 00 at the bifurcation surfaces. The Killing field that is normal to the cosmological event horizon is denoted by $\xi^{a}$. Then the Killing vector $\hat{\xi}^{a}$ which coincides with the generators of the black hole event horizon can be expressed in the form[6]
@@ 4 The non-stationary first law for de Sitter horizons
In this section we utilize the covariant phase space formalism to derive both the non-stationary comparison first law for de Sitter black holes, and the non-stationary physical process first law for cosmological event horizons.
@@ 4.1 The comparison first law
Since for a stationary de Sitter black hole, $J_{H}=\int_{\mathcal{C}_{H}}\boldsymbol{Q}_{g}[\varphi]=\frac{1}{16\pi}\int_{\mathcal{C}_{H}}\boldsymbol{\epsilon}_{ab}\nabla^{a}\varphi^{b}$ recovers the definition of the angular momentum of the black hole[6]. Therefore the non-stationary comparison first law between an arbitrary black hole horizon cross-section $\mathcal{C}_{H}$ and an arbitrary cosmological horizon cross-section $\mathcal{C}_{C}$ reads
@@ 4.2 The physical process first law
Figure 3: The physical process first law for the de Sitter horizon. An external matter source denoted by $\delta T_{ab}$ passes through the cosmological event horizon between two arbitrary cross-sections $\mathcal{C}_{C}(v_{1})$ and $\mathcal{C}_{C}(v_{2})$. The physical process first law computes the entropy change between $\mathcal{C}_{C}(v_{1})$ and $\mathcal{C}_{C}(v_{2})$ in this process.
@@
$\Delta\delta M=-\int_{v_{1}}^{v_{2}}\mathrm{d}v\int_{\mathcal{C}_{C}(v)}\mathrm{d}A\ \delta T_{ab}\xi^{a}k^{b},$ (4.17)
@@ 5 Conclusion and discussion
We have lifted the construction of the dynamical entropy to cosmological event horizons in asymptotically de Sitter spacetimes. By applying the Noether charge method to non-stationary perturbations of a stationary de Sitter black hole spacetime, we have demonstrated that dynamical entropy satisfies both the non-stationary comparison first law and the non-stationary physical process first law. All of these further encourage us to extend the dynamical entropy to more general horizons[22, 23], such as expansion and shear free horizons.